OpenStudy (anonymous):

solve by factoring. 4x^2-5x+1=0

OpenStudy (nathan917):

Simplifying 4x2 + -5x + 1 = 0 Reorder the terms: 1 + -5x + 4x2 = 0 Solving 1 + -5x + 4x2 = 0 Solving for variable 'x'. Factor a trinomial. (1 + -4x)(1 + -1x) = 0 Subproblem 1 Set the factor '(1 + -4x)' equal to zero and attempt to solve: Simplifying 1 + -4x = 0 Solving 1 + -4x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -4x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -4x = 0 + -1 -4x = 0 + -1 Combine like terms: 0 + -1 = -1 -4x = -1 Divide each side by '-4'. x = 0.25 Simplifying x = 0.25 Subproblem 2 Set the factor '(1 + -1x)' equal to zero and attempt to solve: Simplifying 1 + -1x = 0 Solving 1 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -1x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -1x = 0 + -1 -1x = 0 + -1 Combine like terms: 0 + -1 = -1 -1x = -1 Divide each side by '-1'. x = 1 Simplifying x = 1 Solution x = {0.25, 1} pls fan and medal.

OpenStudy (anonymous):

thank you (: